find the measure of an angle if its measure is 5 degrees less than four times the measure of its complement.
Answer by checkley77(12844)

4x-5+x=90
5x=90+5
5x=95
x=95/5
x=19 degrees for the smaller angle.
90-19=71 degrees for the larger angle.
Proof:
4*19-5+19=90
76-5+19=90
90=90